Nested squares in python programming

(f) Nested squares with outer side length of 2*N+1 and each inner square spaced as shown. Note that any two s, each from a
Answer #1

python code:

def nested_squares(N):
n = 2*N+1

# no of rows in square
for i in range(1,n+1):
  
# no of columns in square
for j in range(1,n+1):
  
# print the star in square
if ((i == 1 or i == n or j == 1 or j == n) or
(i >= 3 and i <= n - 2 and j >= 3 and j <= n - 2) and
(i == 3 or i == n - 2 or j == 3 or j == n - 2) or (i== (n+1)/2 and j== (n+1)/2)):
  
print("*",end="") # print star
  
else:
  
print(" ",end="") # print space
  
print()
  
  
# call the function
nested_squares(4)

code and output:

Eile Edit Format Run Options Window def nested squares (N) 2*N+1 n = # no of rows in 3quare for i in range (i,n+1): # no of c

//for any clarification, please do comments. if you found this solution useful, please give me thumbs up.